### v1

@@ 2 General Relativity: assumptions and shortcomings
There is the foundational problem of how we can justify the coincidence between gravitational and inertial mass, which is the basis of the formulation of the WEP. For Newton, the mass of any body, understood as the property of the body itself to respond to a force, corresponded to its ”weight”, which is its property to respond to gravity. In modern terms, we would say that inertial mass $m_{I}$ is equal to passive gravitational mass $m_{G}$, terms coined by Bondi [51]. Einstein said that it was precisely the famous Eötvös experiments on the equivalence between $m_{I}$ and $m_{G}$ that directly inspired him in the formulation of the EEP [18] and which, in fact, constitutes one of its cornerstone.
Today, for many people, this equivalence might seem obvious, but back then, it was not, and it would not be even today if we ”forgot” to acknowledge this principle. In other words, prima facie, there are no reasons to postulate the identity $m_{G}\equiv m_{I}$, and it is not related to some fundamental symmetry. Einstein himself embraced it from empirical reasons. So apart from observations, how one could even imagine this equivalence?
@@
It is important to underline that in GR, torsion of spacetime is set to zero a priori, since with the imposition of the SEP, Einstein chose the symmetric connection, the Levi-Civita one, and so curvature of spacetime. In 1922, Cartan explored a different direction, considering a natural extension of GR constituted not only by the Levi-Civita connection, but also by the torsion tensor, that is the antisymmetric part of a metric compatible affine connection. In this way, he developed a geometric formulation where he suggested that torsion can be physically related to the intrinsic (quantum) angular momentum of matter and it vanishes in vacuum [5]. Einstein himself, in the period 1923 – 1933, tried different geometries for the construction of a unified field theory [30, p.57]. In 1928, he published his first paper on “fernparallelism”, or teleparallelism [52]. As we will see in more detail, since TEGR is found to be dynamically equivalent to GR, classical tests that were understood to confirm the curvature of spacetime can similarly be understood as confirming the torsion of spacetime. This is known as the problem of geometric under-determination [12, 14].
A priori, why prefer curvature over torsion of spacetime? Similar considerations can be applied also to non-metricity, but curvature and torsion are easier to think about metaphysically. It is natural to our mind to think that a massive object can cause curvature of spacetime. However, prima facie, it is natural, in a similar way, to think that a massive object could cause also torsion. Consider, for instance, a rotating black hole.
@@ 3 Equivalent Gravities
$\{M,g_{\mu\nu}\}\,,$ (6)
due to the imposition of the EP. On the contrary, following the Palatini approach [54], the metric $g_{\mu\nu}$ and the connection $\Gamma^{\rho}_{\mu\nu}$ can be varied independently. In this case, spacetime is assigned by the triple:
$\{M,g_{\mu\nu},\Gamma^{\rho}_{\mu\nu}\},$ (7)
@@
Einstein himself recognized as significant the Palatini method, since it represents a simplification of the relativistic formalism [29, p. XXIII], or a generalization, we would say. With the Palatini approach, the connection $\Gamma^{\rho}_{\mu\nu}$ can be written in a more general form considering the affine connection [7, 8]:
@@
Where $\begin{Bmatrix}\rho\\ \mu\nu\end{Bmatrix}$ is the Levi-Civita connection, $K^{\rho}_{\mu\nu}$ and $L^{\rho}_{\mu\nu}$ are the contortion and the disformation tensors, respectively [8]:
@@
$T^{\mu}_{\nu\rho}=\Gamma^{\mu}_{\rho\nu}-\Gamma^{\mu}_{\nu\rho}\neq 0,$ (12)
$Q_{\mu\nu\rho}=\nabla_{\mu}g_{\nu\rho}=\partial_{\mu}g_{\nu\rho}-\Gamma^{\lambda}_{\mu\nu}g_{\mu\rho}-\Gamma^{\lambda}_{\mu\rho}g_{\lambda\nu}\neq 0.$ (13)
@@
As anticipated, several authors [5, 11, 53, 7] claim that TEGR and STEGR can be formulated to be dynamically equivalent to GR in multiple features. Thus, in the following we are going to analyse the features in which the equivalence arises.
@@ 3.1 Equivalence of Lagrangians
where $\mathcal{L}_{GR}=R$ and R is the Ricci curvature scalar; $\mathcal{L}_{m}$ is the matter Lagrangian.
In TEGR and STEGR the same dynamics can be recovered, up to a boundary term. In TEGR, with $\mathcal{L}_{TEGR}=-T$, i.e. the torsion scalar, we have [5]:
@@
$Q=g^{\mu\nu}(L^{\alpha}_{\beta\mu}L^{\beta}_{\nu\alpha}-L^{\alpha}_{\beta\alpha}L^{\beta}_{\mu\nu})=R+\nabla_{\mu}(Q^{\mu}-\bar{Q}^{\mu}),$ (19)
@@
First of all, the equivalence among the three theories is evident at Lagrangian level, up to a boundary term. As mentioned, this equivalence does not hold for extensions like $f(R)$, $f(T)$, and $f(Q)$ [9].
@@ 3.2 Equivalence of the field equations
3.3 The role of the geometric Strong Equivalence Principle
Finally, theorists have found another significant dynamical equivalence, that is the recovery of the SEP [53, 5].
Briefly, TEGR is built upon tetrads $e^{A}_{\mu}$, which describe gravity, and spin connections $\omega^{A}_{B\mu}$, which account for inertial effects. Tetrad fields are geometric constructions which establish a relation between the manifold and its tangent spaces as a soldering agent. On the other hand, the spin connections account for inertial effects in rotated frames. The coordinates of the two frames are related with Lorentz transformations.
One can write also GR with tetrads $\overset{\circ}{\omega}^{A}_{B\mu}$, relating the Levi-Civita connection with the Lorentz connection [7, 5]:
@@ 3.3 The role of the geometric Strong Equivalence Principle
which account for both gravitational and inertial effects in GR. In this way, as shown by several authors, from the General Covariance Principle we can demonstrate the relation between GR and STEGR in the following way [53, 5]:
@@
where $K^{C}_{B\mu}$ accounts for gravitation in TEGR and $\overset{\circ}{\omega}^{C}_{B\mu}$ for both gravitation and inertia in GR. Therefore, in a LIF, where the GR spin connection vanishes $\overset{\circ}{\omega}^{A}_{B\mu}=0$, we obtain the identity between the inertial effects and gravitation in TEGR:
@@
which recover the geometric formulation of the SEP defined in Sec. 2. Here it is evident how in TEGR we have the separation of gravitational and inertial effects, the former identified by the contortion tensor and the latter by the spin connection. For some authors, this possibility of separation is one of the most important properties of TEGR [53].
@@
without curvature and torsion, and by a particular transformation of coordinates, we can arrive at the coincident gauge [55, 5]:
$\Gamma^{\alpha}_{\mu\nu}=\frac{\partial x^{\alpha}}{\partial\xi^{\lambda}}\partial_{\mu}\partial_{\nu}\xi^{\lambda}=0\,.$ (32)
The vanishing of the connection physically means that the origin of the tangent space and the one of the manifold are coincident, which is the geometric formulation of the SEP [5]. The geometric formulation of the SEP is then recovered also in STEGR, via the coincident gauge.
Therefore, as we can see, in TEGR and STEGR, the SEP is not postulated but emerges as a result of a gauge choice.
@@
  • The geometric formulation of SEP is recovered in TEGR and STEGR, even though such a principle is not at their foundation.
@@ 4.1 High degree of under-determination
We have seen that there is a high degree of under-determination among the Geometric Trinity of Gravity, as already pointed out both in the physical [5, 53, 7] and the foundational [12, 13, 14, 56] literature. This causes both various epistemological and metaphysical problems. In this work we focus particularly on if and how we can distinguish experimentally among them. As we will see, there is a crucial hidden difference in the meaning of the EP in the three theories. This fact could represent a difference in the empirical content of the three theories. We think that this difference is strictly related to the assumptions of GR on the spacetime structure and on the fundamental objects of the theory.
@@
These are some of the reasons why some philosophers of physics have delved into this idea and formulated theories of ”confirmation” that make the corresponding intuition more rigorous. This approach (see for instance [57, 58, 59]) exploits Bayesian Confirmation Theory and deeply relies on the practice of exploring and constraining the theory space. In fact, when scientists find that despite substantial efforts, no alternative viable hypothesis are capable of explaining some scientific problem, they tend to place more trust in the existing theory. This is indeed called the ”No Alternatives Argument”. These approaches are called ”meta-empirical”, since they are not empirical in the common sense, but still involve observations. For a general Bayesian formalization of this argument and a proof that it counts as evidence, see Ref. [60], as well as for its limits.
Note that we have not to confuse the use of the concept of non-empirical ”confirmation” as the confirmation of a theory in the traditional sense.
@@
Predictions($Th$) = {$P_{1}$, $P_{2}$, $P_{3}$…}
Predictions($Th^{\prime}$) = {$P_{1}$, $P_{2}$, $\neg P_{3}$}
@@
Predictions($Th$) = {$P_{1}$, $P_{2}$, $P_{3}$…}
Predictions($Th^{\prime}$) = {$P_{1}$, $P_{2}$, $\neg P_{3}$}
Predictions($Th^{\prime\prime}$) = {$P_{1}$, $P_{2}$, $\neg P_{3}$}
@@
Therefore, if this argument is right, this counts as an evidence, a meta-empirical evidence, against the hypothesis $P_{5}$. This counts not as an $empirical$ observation, but as a $meta$-empirical observation. Counting as an observation, it affects the posterior probability of the validity of $P_{5}$, as other empirical evidence [60]. Again, if also all the other competing theories would have had $P_{5}$ as a prediction, this would have been counted as a meta-empirical evidence in its favor. To be clear, in no way this is a posterior evidence with the same strength of an empirical one, but it still counts as evidence.
@@ 4.2 The debate on the Equivalence Principle
As we have seen, despite the fact that TEGR and STEGR are built on different foundation principles with respect to GR, physicists claim that they both recover the SEP, which is of course a necessary condition for a consistent theory of gravity, at least at classical level. However, it is often overlooked that there is a crucial difference in their relation with the SEP. In GR, EP is the fundamental assumption of the theory, while, in TEGR and STEGR, it is not postulated a priori but it is recovered a posteriori. As it will be argued here, this fact could represent a possible important difference in the empirical content of theories in Geometric Trinity, since if it is not fundamental, there is the possibility that, at some level, it could be not valid.
Moreover, this result seems sufficient to regard TEGR and STEGR as different proper theories, instead of mere mathematical reformulations of GR, as some authors suggested [56]. According to that perspective, it was possible that the under-determination among the Geometric Trinity would later turn out to be an ill-posed problem. On the contrary, the recent theoretical advancements seem to have clarified this point.
@@
Pred.($GR$)       = {$L$, $FE$, $S$, $C$, $FEP$}
Pred.($TEGR$)   = {$L$, $FE$, $S$, $C$, $\neg FEP$}
Pred.($STEGR$) = {$L$, $FE$, $S$, $C$, $\neg FEP$}
@@
Finally, $FEP$ stands for Fundamental Equivalence Principle. As demonstrated before, in TEGR and STEGR, the EP is not fundamental, and so we can write $\neg FEP$. In other words, we can say that TEGR and STEGR predict an EMergent Equivalence Principle, or $EMEP$, instead of a $FEP$. This does not mean a prediction for a violation of the EP, but only that TEGR and STEGR do not have the $EP$ at their foundation, i.e. they do not share the prediction $FEP$. In fact, the prediction $EMEP$ is, in some sense, shared by all the three theories, since an $EMEP$ is contained in $FEP$. If the EP is fundamental, it has to be always valid at any level, and so it is also valid at emergent levels, but the contrary is not necessarily the case. In other words, the $FEP$ implies the $EMEP$, but the $EMEP$ does not imply the $FEP$:
@@
Figure 3: The figure shows how the three theories differ with respect to the EP. In GR, EP is at the foundation of the theory and without it the other predictions are not possible; without the root, the branches would not exist. In TEGR and STEGR, on the contrary, the EP is not at the foundation, but it is a lateral branch (the SEP, in particular), recoverable through the general covariance principle and via the coincident gauge, respectively. $L$: Lagrangians, $FE$: field equations, $S$: solutions of the $FE$, $C$: cosmological applications.
As argued, this situation decreases our confidence in that hypothesis on which the equivalent theories do not agree. That is to say, if the argument is correct, our confidence in the fundamentality of the EP is decreased.
This argument is also independent of the history of the theory itself. In fact, imagine that both TEGR and STEGR would be built only with the EP as a foundation; imagine that $L$, $FE$, $S$ and $C$ were derivable only if $FEP$ was valid, i.e. only assuming the EP. Then our confidence in the EP as a necessary principle of any consistent theory of gravity would be increased. In the terms previously defined, this observation would have counted as a meta-empirical evidence for the hypothesis $FEP$, and consequently for a metric theory of gravity. It would have been considered evidence in favor of $FEP$ because, after searching for an alternative hypothesis to $FEP$ to build a viable and coherent theory of gravity, physicists would not have found it. Therefore, as shown above, the posterior probability of the $FEP$ hypothesis would be increased. But the results are pointing to exactly the opposite. In fact, both TEGR and STEGR can recover the EP, but it is not at their foundation, i.e. they do not predict $FEP$. See Fig. 3.
More precisely, any viable theory of gravity has to recover the EP, at least at the scales and at the levels of accuracy of the present experiments. So the fact that TEGR and STEGR can recover the EP is an important feature. This is one of the reason why it is difficult to distinguish empirically among the three theories, since TEGR and STEGR necessarily have to recover the experimental tests of GR. But properly the fact that the EP is not at the foundation of them constitutes a possible important difference from the empirical content.
In other words, this argument shows that EP is not a principle necessary present in all viable theories of gravity, but it is an assumption on which GR is built.

### v2

@@ 2 General Relativity: assumptions and shortcomings
There is the foundational problem of how we can justify the coincidence between gravitational and inertial mass, which is the basis of the formulation of the WEP. For Newton, the mass of any body, understood as the property of the body itself to respond to a force, corresponded to its “weight” , which is its property to respond to gravity. In modern terms, we would say that inertial mass $m_{I}$ is equal to passive gravitational mass $m_{G}$, terms coined by Bondi [60]. Einstein said that it was precisely the famous Eötvös experiments on the equivalence between $m_{I}$ and $m_{G}$ that directly inspired him in the formulation of the EEP [21] and which, in fact, constitutes one of its cornerstone.
Today, for many people, this equivalence might seem obvious, but back then, it was not, and it would not be even today if we “forgot” to acknowledge this principle. In other words, prima facie, there are no reasons to postulate the identity $m_{G}\equiv m_{I}$, and it is not related to some fundamental symmetry. Einstein himself embraced it from empirical reasons. So apart from observations, how one could even imagine this equivalence?
@@
It is important to underline that in GR, torsion of spacetime is set to zero a priori, since with the imposition of the SEP, Einstein chose the symmetric connection, the Levi-Civita one, and so curvature of spacetime. In 1922, Cartan explored a different direction, considering a natural extension of GR constituted not only by the Levi-Civita connection, but also by the torsion tensor, that is the antisymmetric part of a metric compatible affine connection. In this way, he developed a geometric formulation where he suggested that torsion can be physically related to the intrinsic (quantum) angular momentum of matter and it vanishes in vacuum [5]. Einstein himself, in the period 1923 – 1933, tried different geometries for the construction of a unified field theory [34, p. 57]. In 1928, he published his first paper on “fernparallelism”, or teleparallelism [61]. As we will see in more detail, since TEGR is found to be dynamically equivalent to GR, classical tests that were understood to confirm the curvature of spacetime can similarly be understood as confirming the torsion of spacetime. This is known as the problem of geometric under-determination [12, 14].
A priori, why prefer curvature over torsion of spacetime? It seems natural to our minds to think of a massive object as causing curvature of spacetime. However, prima facie, it is natural, in a similar way, to think that a massive object could also cause torsion. Consider, for instance, a rotating black hole. Therefore, curvature of spacetime is not, a priori, more probable than torsion. Similar considerations can be applied also to non-metricity, but curvature and torsion are intuitively easier to think about metaphysically.
@@ 3 Equivalent Gravities
$\langle M,g_{\mu\nu}\rangle\,$, (5)
due to the imposition of the SEP. On the contrary, following the Palatini approach [63], the metric $g_{\mu\nu}$ and the connection $\Gamma^{\rho}_{\mu\nu}$ can be varied independently. In this case, spacetime is assigned by the triple:
$\langle M,g_{\mu\nu},\Gamma^{\rho}_{\mu\nu}\rangle,$ (6)
@@
Einstein himself recognized as significant the Palatini method, since it represents a simplification of the relativistic formalism [33, p. XXIII], or a generalization, we would say. With the Palatini approach, the connection $\Gamma^{\rho}_{\mu\nu}$ can be written in a more general form considering the affine connection [7, 8]:
@@
Where $\begin{Bmatrix}\rho\\ \mu\nu\end{Bmatrix}$ is the Levi-Civita connection, $K^{\rho}_{\mu\nu}$ and $L^{\rho}_{\mu\nu}$ are the contortion and the distortion tensors, respectively [8]:
@@
$T^{\mu}_{\nu\rho}=\Gamma^{\mu}_{\rho\nu}-\Gamma^{\mu}_{\nu\rho},$ (11)
$Q_{\mu\nu\rho}=\nabla_{\mu}g_{\nu\rho}=\partial_{\mu}g_{\nu\rho}-\Gamma^{\lambda}_{\mu\nu}g_{\mu\rho}-\Gamma^{\lambda}_{\mu\rho}g_{\lambda\nu}.$ (12)
@@
As anticipated, several authors [5, 11, 62, 7] claim that TEGR and STEGR can be formulated to be equivalent to GR in multiple features. Thus, in the following we are going to analyse the features in which the equivalence arises.
@@ 3.1 Equivalence of Lagrangians
where $\mathcal{L}_{GR}=\mathcal{R}$ and $\mathcal{R}$ is the Ricci curvature scalar; $\mathcal{L}_{m}$ is the matter Lagrangian. In TEGR and STEGR the same dynamics can be recovered, up to a boundary term.
Teleparallel gravity can be built in different ways (see [56] for a recent review). A useful approach to formulating it as a gauge theory [56, p. 16] is to work with tetrads $e^{A}_{\mu}$, which describe gravity, and spin connections $\omega^{A}_{B\mu}$, which account for inertial effects. We follow this tradition [62, 5, 64], not least because it provides an elegant formulation of the SEP. Tetrad fields are geometric constructions which establish a relation between the manifold and its tangent spaces as a soldering agent. On the other hand, the spin connections account for inertial effects in rotated frames. The coordinates of the two frames are related with Lorentz transformations.
Thus, in TEGR, with $\mathcal{L}_{TEGR}=-T$, i.e. the torsion scalar, we have [5]:
@@
$Q=g^{\mu\nu}(L^{\alpha}_{\beta\mu}L^{\beta}_{\nu\alpha}-L^{\alpha}_{\beta\alpha}L^{\beta}_{\mu\nu})=\mathcal{R}+\nabla_{\mu}(Q^{\mu}-\bar{Q}^{\mu}),$ (18)
@@
First of all, the equivalence among the three theories is evident at Lagrangian level. In fact, the actions of the three theories differ by boundary terms, which are dynamically passive, since the spacetime is asymptotically flat at infinity and thus these terms give no contribution [5]. As mentioned, this equivalence does not hold for extensions like $f(\mathcal{R})$, $f(T)$, and $f(Q)$ [9].
@@ 3.2 Equivalence of the field equations
3.3 The role of the Strong Equivalence Principle
Finally, theorists have found another significant empirical equivalence, that is the recovery of the SEP [62, 5].
As for TEGR, one can also formulate GR in the tetrad formalism, where the Levi-Civita connection is translated into a Lorentz connection [7, 5]:
@@ 3.3 The role of the Strong Equivalence Principle
where $\overset{\circ}{\omega}^{A}_{B\mu}$ accounts for both gravitational and inertial effects in GR. From the General Covariance Principle in its active formulation, we can obtain the coupling gravitational prescription. This principle states that the laws of physics - initially written in a special-relativistic context - must retain their form also in the presence of gravity, requiring invariance under spacetime diffeomorphisms. In the case of GR, for a general field $\Psi$ [62, 5], it is:
$\partial_{\mu}\Psi\quad\longrightarrow D_{\mu}\Psi=\partial_{\mu}\Psi+\frac{1}{2}\overset{\circ}{\omega}{}^{C}_{\ B\mu}S^{B}_{\ C}\Psi$ (28)
where $S^{B}_{\ C}$ are the Lorentz generators.
The analogous coupling in TEGR can be similarly derived. Transitioning from the special-relativistic to the case with gravity, incorporating the appropriate contortion tensor in the Lorentz covariant derivative, we obtain:
$\partial_{\mu}\Psi\quad\longrightarrow D_{\mu}\Psi=\partial_{\mu}\Psi+\frac{1}{2}\left(\omega^{C}_{\ B\mu}-K^{C}_{\ B\mu}\right)S^{B}_{\ C}\Psi$ (29)
The combination of (28) and (29) yields the relation between the gravitational and inertial effects in TEGR and GR [62, 5]:
@@
where $\omega^{C}_{B\mu}$ and $K^{C}_{B\mu}$ account respectively for inertia and gravitation in TEGR, and ${\omega}^{C}_{B\mu}$ for both gravitation and inertia in GR. Therefore, choosing a LIF, where the GR spin connection vanishes $\overset{\circ}{\omega}^{A}_{B\mu}=0$, we obtain the identity between the inertial effects and gravitation in TEGR [62, 5]:
@@
which recover the second formulation of SEP defined in Sec. 2, SEP2,¬EEP, which does not assume the EEP, or in the words by Read and Teh, it does not conceptually unify gravity and inertia [24, p. 3486]. In fact, here it is evident how in TEGR we have the separation of gravitational and inertial effects, the former identified by the contortion tensor and the latter by the spin connection. For some authors, this possibility of separation is one of the most important properties of TEGR [62].
For a recovery of SEP1,EEP in the TEGR context, see [24, p. 3491].
@@
without curvature and torsion, and by a particular transformation of coordinates in a point $p$, we can arrive at the coincident gauge [65, 5, 59]:
$\Gamma^{\alpha}_{\mu\nu}(p)=\frac{\partial x^{\alpha}}{\partial\xi^{\lambda}}\partial_{\mu}\partial_{\nu}\xi^{\lambda}=0\,.$ (33)
The vanishing of the connection physically means that the origin of the tangent flat space and the one of the manifold are coincident [5], which is the formulation of the SEP defined in Sec. 2. The SEP is then recovered also in STEGR, via the coincident gauge.
Recovery of been considered one of the most important ways in which TEGR and STEGR demonstrate their empirical equivalence with GR [5, 62]. Note also that this empirical equivalence is distinct from the equivalence given by the dynamics, i.e. by the action (Sec. 3.1). In fact, as Wolf &\& Read show, the empirical content of a theory is not exhausted by its dynamics [19]. The recovery of the SEP in the explored regimes and at the tested levels is, of course, essential for any theory of gravity to be considered consistent, given the amount of high-precision experimental tests (see Sec. 5.2 for the state of the art). Therefore, these results are highly significant. However, in TEGR and STEGR, the SEP is not postulated as fundamental, as it is in GR, but instead emerges as the result of a possible gauge choice. This structural difference, as we will argue in the following sections, could imply differences in empirical content.
@@
  • The SEP is recovered in TEGR and STEGR, even though such a principle is not at their foundation.
@@ 4.1 High degree of under-determination
We have seen that there is a high degree of under-determination among the Geometric Trinity of Gravity, as already pointed out both in the physical [5, 62, 7] and the foundational [12, 13, 14, 17] literature. This causes both various epistemological and metaphysical problems. In this work we focus particularly on if and how we can distinguish experimentally among them. Although the three theories are dynamically equivalent, and the recovery of SEP is often regarded as one of the key features through which TEGR and STEGR demonstrate their empirical equivalence with GR, we argue that the differing foundational status of SEP across the three theories may nonetheless lead to differences in empirical content. We think that this difference is strictly related to the assumptions of GR on the spacetime structure and on the fundamental objects of the theory.
@@
These are some of the reasons why some philosophers of physics have delved into this idea and formulated theories of “confirmation” that make the corresponding intuition more rigorous. This approach (see for instance [66, 67, 68]) exploits Bayesian Confirmation Theory and deeply relies on the practice of exploring and constraining the theory space. In fact, when scientists find that despite substantial efforts, no alternative viable hypothesis are capable of explaining some scientific problem, they tend to place more trust in the existing theory. This is indeed called the “No Alternatives Argument”. These approaches are called “meta-empirical”, since they are not empirical in the common sense, but still involve observations. For a general Bayesian formalization of this argument and a proof that it counts as evidence, see Ref. [69], as well as for its limits.
Note that we have not to confuse the use of the concept of non-empirical “confirmation” as the confirmation of a theory in the traditional sense.
@@
Predictions(ThTh) = $P_{1}$, $P_{2}$, $P_{3}$…}
Predictions($Th^{\prime}$) = $P_{1}$, $P_{2}$, $\neg P_{3}$ …}
@@
Predictions(ThTh) = $P_{1}$, $P_{2}$, $P_{3}$…}
Predictions($Th^{\prime}$) = $P_{1}$, $P_{2}$, $\neg P_{3}$ …}
Predictions($Th^{\prime\prime}$) = $P_{1}$, $P_{2}$, $\neg P_{3}$ …}
@@
Therefore, if this argument is right, this counts as an evidence, a meta-empirical evidence, against the hypothesis $P_{5}$. This counts not as an empiricalempirical observation, but as a metameta-empirical observation. Counting as an observation, it affects the posterior probability of the validity of $P_{5}$, as other empirical evidence [69]. Again, if also all the other competing theories would have had $P_{5}$ as a prediction, this would have been counted as a meta-empirical evidence in its favor. To be clear, in no way this is a posterior evidence with the same strength of an empirical one, but it still counts as evidence.
@@ 4.2 The debate on the Equivalence Principle
As we have seen, despite the fact that TEGR and STEGR are built on different foundation principles with respect to GR, physicists claim that they both recover the SEP, which is of course a necessary condition for a consistent theory of gravity, at least at classical level. However, it is often overlooked that there is a crucial difference in their relation with the SEP. In GR, SEP is a fundamental assumption of the theory, while, in TEGR and STEGR, it is not postulated a priori but it is recovered a posteriori. As it will be argued here, this structural difference could represent a possible important distinction in the empirical content of theories in Geometric Trinity, since if it is not fundamental, there is the possibility that, at some level, it could be not valid.
Moreover, this result seems sufficient to regard TEGR and STEGR as different proper theories, instead of mere mathematical reformulations of GR, as some authors have suggested based on their dynamical equivalence [17]. According to that perspective, it was possible that the under-determination among the Geometric Trinity would later turn out to be an ill-posed problem. In contrast, recent theoretical advancements seem to have clarified this point. This is consistent with other works that acknowledge the lack of categorical equivalence between GR and TEGR [71, 18, 59, 56].
@@
Pred.(GRGR)       = {$L$, FEFE, $S$, $C$, FEPFEP}
Pred.(TEGRTEGR)   = {$L$, FEFE, $S$, $C$, $\neg FEP$}
Pred.(STEGRSTEGR) = {$L$, FEFE, $S$, $C$, $\neg FEP$}
@@
Finally, $FEP$ stands for Fundamental Equivalence Principle. As demonstrated before, in TEGR and STEGR, the SEP is not fundamental, and so we can write $\neg FEP$. In other words, we can say that TEGR and STEGR predict an EMergent Equivalence Principle, or $EMEP$, instead of a $FEP$. This does not mean a prediction for a violation of the SEP, but only that TEGR and STEGR do not have the SEP at their foundation, i.e. they do not share the prediction $FEP$. In fact, the prediction $EMEP$ is, in some sense, shared by all the three theories, since an $EMEP$ is contained in $FEP$. If the SEP is fundamental, it has to be always valid at any level, and so it is also valid at emergent levels, but the contrary is not necessarily the case. In other words, the $FEP$ implies the $EMEP$, but the $EMEP$ does not imply the $FEP$:
@@
Figure 3: The figure shows how the three theories differ with respect to the EP (the SEP, in particular). In GR, EP is at the foundation of the theory and without it the other predictions are not possible; without the root, the branches would not exist. In TEGR and STEGR, on the contrary, the EP is not at the foundation, but it is a lateral branch, recoverable through the general covariance principle and via the coincident gauge, respectively. $L$: Lagrangians, FEFE: field equations, $S$: solutions of the $FE$, $C$: cosmological applications.
As argued, this situation decreases our confidence in that hypothesis on which the equivalent theories do not agree. That is to say, if the argument is correct, our confidence in the fundamentality of the SEP is decreased.
This argument is also independent of the history of the theory itself. In fact, imagine that both TEGR and STEGR would be built only with the SEP as a foundation; imagine that $L$, FEFE, $S$ and $C$ were derivable only if $FEP$ was valid, i.e. only assuming the SEP. Then our confidence in the SEP as a necessary principle of any consistent theory of gravity would be increased. In the terms previously defined, this observation would have counted as a meta-empirical evidence for the hypothesis $FEP$, and consequently for a metric theory of gravity. It would have been considered evidence in favor of $FEP$ because, after searching for an alternative hypothesis to $FEP$ to build a viable and coherent theory of gravity, physicists would not have found it. Therefore, as shown above, the posterior probability of the $FEP$ hypothesis would be increased. But the results are pointing to exactly the opposite. In fact, both TEGR and STEGR can recover the SEP, but it is not at their foundation, i.e. they do not predict $FEP$. See Fig. 3.
More precisely, any viable theory of gravity has to recover the SEP, at least at the scales and at the levels of accuracy of the present experiments. So the fact that TEGR and STEGR can recover the SEP is an important feature. This is one of the reason why it is difficult to distinguish empirically among the three theories, since TEGR and STEGR necessarily have to recover the experimental tests of GR. But properly the fact that the SEP is not at the foundation of them constitutes a possible important difference from the empirical content.
In other words, this argument shows that SEP is not a principle necessary present in all viable theories of gravity, but it is an assumption on which GR is built.